OpenStudy (anonymous):

Solve -2 ( 3- m )+ 15 = 6m - 4 ( m - 20 ) Please show all of your work

OpenStudy (anonymous):

ok so use PEMDAS for this one. distribute -2 on left side to 3-m: -6+m+15=6m-4(m-20) 9+m=6m-4m+80 Distribute the -4 and add -6 and 15. 9+m=2m+80 I think you can solve from here.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

np.

OpenStudy (anonymous):

it cant add up to 80

OpenStudy (anonymous):

What do you mean? you need to solve for m.

OpenStudy (anonymous):

i got 9 = 80

OpenStudy (anonymous):

No you need to find m. So: 9+m=2m+80 9-m=80 Subtract 2m from both sides. -m=71 Subtract 9 from both sides m=-71 Divide by -1.

OpenStudy (anonymous):

-71divide by -1 = 71 so m= 71

OpenStudy (anonymous):

no. -m=positive 71. So 71/-1 is -71.

OpenStudy (anonymous):

oh ok so the - become +

OpenStudy (anonymous):

yea.